### v2

# Interior curvature estimate for curvature quotient equations on convex hypersurfaces

Abstract.
We study interior curvature estimates for convex graphs which satisfy the quotient equation $\frac{\sigma_{n}}{\sigma_{n-2}}(\lambda)=f(X)>0$ in this paper.
Key words and phrases:
Hessian quotient equation, interior curvature estimate, fully nonlinear elliptic equations
2020 Mathematics Subject Classification:
Primary 35J60, 35B45; Secondary 53C21
1. Introduction
Interior curvature estimates play a fundamental role in understanding the regularity of solutions to geometric partial differential equations. In this paper, we focus on deriving interior curvature bounds for convex solutions to a specific class of curvature equations expressed as
(1.1) $\frac{\sigma_{n}}{\sigma_{n-2}}(\lambda(X))=f(X),\quad f(X)>0,$
where $\sigma_{k}$ represents the $k$-th elementary symmetric polynomial, and $X,\lambda(X)$ represent the position vector of $M$ and the principal curvatures of hypersurfaces $M\subset\mathbb{R}^{n+1}$ at $X$ respectively.
For equation (1.1), we establish the following interior $C^{2}$ estimate.
Theorem 1.1.
For $n\geq 3$, suppose that $M=(x,u(x))$ is a local $C^{4}$ graph over a ball $x\in B_{r}\subset\mathbb{R}^{n}$ with positive principal curvatures $\lambda=(\lambda_{1},\cdots,\lambda_{n})>0$ and that it is a solution to equation (1.1) in $B_{r}$. And $f\in C^{2}(B_{r})$ is a positive function. Then there is a positive constant $C$ such that
$\text{sup}_{B_{\frac{r}{2}}}|\lambda_{i}|\leq C,$
where $C$ depends only on $n,r,\|f\|_{C^{2}(B_{r})},inf_{B_{r}}|f|,\|M\|_{C^{1}(B_{r})}$.
Equation (1.1) is a special case of the equation (1.2) under conditions (1.41.7). We assume that $F\in C^{2}(\Gamma)\cap C^{0}(\bar{\Gamma})$ is a symmetric function defined on a bounded, convex and symmetric region $\Gamma\subset\mathbb{R}^{n},\Gamma\neq\mathbb{R}^{n}$, with $0\in\partial\Gamma$ and such that $\Gamma+\Gamma_{+}\subset\Gamma$, where $\Gamma_{+}$ is the positive cone in $\mathbb{R}^{n}$ and $u$ is a function defined in $\Gamma$.
At any point $x\in\Gamma$, the principal curvatures $\lambda=(\lambda_{1},\cdots,\lambda_{n})$ of the graph $X=(x,u(x))$ satisfy curvature equations
(1.2) $F[\lambda(u)]=f(X,\nu(X)),$
where $\nu(X)$ is the outer normal vector at $X$. Similarly, the eigenvalue of hessian matrix $D^{2}u$ satisfy k-Hessian equations
(1.3) $F[\lambda(D^{2}u)]=g(x,u,Du),$
In the seminal work [2] and [3], L. Caffarelli, L. Nirenberg, and J. Spruck studies a wide class of operators $\Phi$, which can be represented by a smooth symmetric function $\Phi(D^{2}u)=F(\lambda(D^{2}u)))$, where $\lambda$ are the eigenvalues of $D^{2}u$. $F$ is defined in an open convex cone $\Gamma\subset\mathbb{R}^{n}$ with vertex at the origin and containing the positive cone $\lambda\in\mathbb{R}^{n}|\forall i,\lambda_{i}>0$} such that
(1.4) $F>0\hskip 8.53581ptin\hskip 8.53581pt\Gamma,$ $\hskip 14.22636ptF=0\hskip 8.53581pton\hskip 8.53581pt\partial\Gamma,$
[…]
(1.6) $\sum_{i=1}^{n}F^{ii}$ $\geq\omega_{1}\hskip 8.53581pton\hskip 8.53581pt\Gamma_{\mu_{1},\mu_{2}},$
(1.7) $\sum_{i=1}^{n}F^{ii}\lambda_{i}$ $\geq\omega_{2}\hskip 8.53581pton\hskip 8.53581pt\Gamma_{\mu_{1},\mu_{2}},$
where $\Gamma_{\mu_{1},\mu_{2}}=\{\lambda\in\Gamma:\mu_{1}\leq F[u]\leq\mu_{2}\}$ for any $\mu_{2}\geq\mu_{1}\geq 0$, and $\omega_{1},\omega_{2}$ are positive constants depending on $\mu_{1},\mu_{2}$, and $F^{ii}=\frac{\partial F}{\partial\lambda_{i}}$.
As notable examples of this class, k-Hessian operators ($\sigma_{k}^{1/k}$) and curvature quotient operators $\left(\frac{\sigma_{k}}{\sigma_{l}}\right)^{1/(k-l)}$, where $n\geq k>l\geq 1$, are important research objects in geometric analysis due to their close connection with curvatures. For instance, $\sigma_{k}$ is corresponding to the mean curvature, scalar curvature, and Gauss curvature when $k=1,2,n$, respectively.
The study of interior curvature and $C^{2}$ estimates for solutions to equations (1.2) and (1.3) has been an area of extensive and continuous research. This topic is of significant interest as it allows for the investigation of solution regularity in the absence of boundary data. This problem could be dated back to [9] by Heinz in 1959, in which he established an interior $C^{2}$ estimate for the Monge–Ampère equation — det($D^{2}u$)=$f(x,u,Du)$ in $\mathbb{R}^{2}$. Subsequently, similar results were obtained by Chen, Han, and Ou in [4], as well as Liu in [11], using different methods.
However, Pogorelov demonstrated that the interior $C^{2}$ estimate fails for the Monge –Ampère equation in dimensions $n\geq 3$, as shown by a counterexample in [16]. Urbas extended this counterexample to conclude that the interior $C^{2}$ estimate fails for the Hessian equation $\sigma_{k}(D^{2}u)=f$ for $k\geq 3$ in [22]. Despite these findings, the case of the Hessian equation $\sigma_{2}(D^{2}u)=f$ remains unresolved, and its general dimensional case is still an open problem.
Notably, there have been significant advancements in addressing this issue under certain restrictions for $n\geq 3$. When $n=3$, Warren and Yuan made a breakthrough by establishing an interior $C^{2}$ estimate for $\sigma_{2}(D^{2}u)=1$ in [24]. Building on this, Qiu generalized the coefficient $1$ in right hand side to $\sigma_{2}(D^{2}u)=f(x,u,Du)$ in [17], and further extended it to the curvature equation $\sigma_{2}(\lambda)=f(X,\nu(X))$ in [18]. For $n=4$, Shankar and Yuan provided an interior estimate for $\sigma_{2}(D^{2}u)=1$ in their latest work [20]. For general dimensions $n$, Mooney addressed the interior estimate problem for strictly convex solutions to $\sigma_{2}(D^{2}u)=1$ in [15]. McGonagle, Song, and Yuan [14] further generalized these results to solutions satisfying $D^{2}u>-c(n)$. In addition, Shankar and Yuan revisited these estimates by modifying the conditions to $D^{2}u>-KI$ in [19] and $D^{2}u>-c(n)\Delta u$ in [20], respectively. Generalizing the right-hand side from $1$ to $f$, Guan and Qiu [6] determined the interior $C^{2}$ estimate for equations form $\sigma_{2}(D^{2}u)=f(x,u,Du)$ and $\sigma_{2}(\lambda)=f(X,\nu(X))$, provided that $\sigma_{3}(\lambda)>-AI$.
Meanwhile, research on Hessian quotient equations has progressed more slowly due to their complicated structure. For $1\leq k\leq n$, Sheng, Urbas, and Wang [21] and Guan and Zhang [8] discussed the curvature equations $\frac{\sigma_{k}}{\sigma_{k-1}}(D^{2}u)=f(x,u),\frac{\sigma_{k}}{\sigma_{k-1}}(\lambda)=f(X)$ and the corresponding interior $C^{2}$ estimate for admissible solutions. Nevertheless, in dimensions $3$ and $4$, Chen, Warren, and Yuan [5], as well as Wang and Yuan [23], successfully extended the interior $C^{2}$ estimate for convex solutions by employing special Lagrangian structures of the equations $\frac{\sigma_{3}}{\sigma_{1}}(D^{2}u)=1.$ Zhou [25] enhanced the results by incorporating a twisted Lagrangian structure into the equation $\frac{\sigma_{3}}{\sigma_{1}}(D^{2}u)=f(x,u)$. The three-dimensional case was also achieved by Lu [12] by introducing the Legendre transform. Recently, a significant breakthrough was achieved in [13], where Lu improved the interior $C^{2}$ estimates for convex solutions to equations of the same form to higher dimensions $n$. He also showed that these estimates fail for any $\frac{\sigma_{k}}{\sigma_{l}}(D^{2}u)=f$ when $1\leq l<k\leq n$ and $k-l\geq 3$, following Urbas’ approach in [22]. The remaining case for Hessian quotient equations $\frac{\sigma_{k}}{\sigma_{k-2}}(D^{2}u)=f$ when $3\leq k<n$ and curvature quotient equations $\frac{\sigma_{k}}{\sigma_{l}}(\lambda)=f$ where $1\leq l<k\leq n$ are still unknown.
Inspired by Lu’s work [13], we revisit the Jacobi inequality on Riemannian manifolds. In contrast to the integral method developed by Lu, which draws on the framework of Shankar and Yuan [20], our approach is pointwise, thus avoiding the technical difficulties of applying the Legendre transform on Riemannian manifolds. Building on the auxiliary function technique of Guan and Qiu [6], we develop a new case-by-case analysis tailored to the structure of Hessian quotient equations and obtain a structurally similar result. The cases $n=3$ or $4$ have been well studied, while higher dimensions remain less understood. Our method applies to general $n\geq 3$ and may offer a modest contribution.
The organization of the paper is as follows. We collect and prove some facts related to the Hessian quotient operator $\frac{\sigma_{n}}{\sigma_{n-2}}$ and state a crucial concavity inequality in Section 2. Then we derive the Jacobi inequality for this Hessian quotient operator on Riemannian manifolds in Section 3. Finally, we finish the proof of Theorem 1.1 in Section 4 with all the facts found in Section 2, 3.
2. Preliminaries
Define the operator $F$ as
$F=\frac{\sigma_{n}}{\sigma_{n-2}}.$
Let $h=(h_{ij})$ denote the second fundamental form of $M$, whose eigenvalues are denoted by $\lambda=(\lambda_{1},\dots,\lambda_{n})\in\mathbb{R}^{n}$. We introduce the following notation
$(\lambda|i)=(\lambda_{1},\dots,\lambda_{i-1},\lambda_{i+1},\dots,\lambda_{n})\in\mathbb{R}^{n-1},$
where $(\lambda|i)$ corresponds to the vector formed by removing the $i$-th component of $\lambda$. Analogously, $(\lambda|ij)$ denotes the vector obtained by deleting both the $i$-th and $j$-th components of $\lambda$.
We first collect some basic formulas for the Hessian operator (see, for instance, [10]).
Lemma 2.1.
For any $\lambda=(\lambda_{1},\cdots,\lambda_{n})\in\mathbb{R}^{n}$, the following identities hold:
$\sigma_{k}(\lambda)=\lambda_{i}\sigma_{k-1}(\lambda|i)+\sigma_{k}(\lambda|i),\quad\sum_{i}\sigma_{k}(\lambda|i)$ $=(n-k)\sigma_{k}(\lambda),$
$\sum_{i}\lambda_{i}\sigma_{k-1}(\lambda|i)=k\sigma_{k}(\lambda),\quad\sum_{i}\lambda_{i}^{2}\sigma_{k-1}(\lambda|i)=\sigma_{1}(\lambda)$ $\sigma_{k}(\lambda)-(k+1)\sigma_{k+1}(\lambda).$
We now summarize fundamental properties of the operator $F$. Proofs of Lemma 2.2 was provided in [13].
Lemma 2.2.
Let $1\leq k<n$ and let $n\geq 2$. Assume the coordinates are rotated such that $h$ is diagonalized at $X(x_{0})$ for $x_{0}\in B_{r}$. Then at $X(x_{0})$,
$F^{pq}=\frac{\partial F}{\partial h_{pq}}=\frac{\sigma_{n}^{pq}}{\sigma_{k}}-\frac{\sigma_{n}\sigma_{k}^{pq}}{\sigma_{k}^{2}},$
$F^{pq,rs}=\frac{\partial^{2}F}{\partial h_{pq}\partial h_{rs}}=\begin{cases}-2\frac{\sigma_{n}^{pp}\sigma_{k}^{pp}}{\sigma_{k}^{2}}+2\frac{\sigma_{n}\left(\sigma_{k}^{pp}\right)^{2}}{\sigma_{k}^{3}},\quad&p=q=r=s,\\ \frac{\sigma_{n}^{pp,rr}}{\sigma_{k}}-\frac{\sigma_{n}^{pp}\sigma_{k}^{rr}}{\sigma_{k}^{2}}-\frac{\sigma_{n}^{rr}\sigma_{k}^{pp}}{\sigma_{k}^{2}}-\frac{\sigma_{n}\sigma_{k}^{pp,rr}}{\sigma_{k}^{2}}+2\frac{\sigma_{n}\sigma_{k}^{pp}\sigma_{k}^{rr}}{\sigma_{k}^{3}},&p=q,r=s,p\neq r,\\ \frac{\sigma_{n}^{pq,qp}}{\sigma_{k}}-\frac{\sigma_{n}\sigma_{k}^{pq,qp}}{\sigma_{k}^{2}},&p=s,q=r,p\neq q,\\ 0,&\textit{otherwise}.\end{cases}$
Moreover, for all $p\neq q$, the inequality $-F^{pq,qp}>0$ holds, and
$-F^{pq,qp}=\frac{F^{pp}-F^{qq}}{\lambda_{q}-\lambda_{p}},\quad\lambda_{p}\neq\lambda_{q},$
where $\lambda_{p}$ and $\lambda_{q}$ denote the eigenvalues of $h$.
Lemma 2.3.
Let $n\geq 3$. Suppose $M$ is a convex graph over $B_{r}\subset\mathbb{R}^{n}$ with coordinates rotated to diagonalize $h$ at $X(x_{0})$, and the eigenvalues arranged as $\lambda_{1}\geq\cdots\geq\lambda_{n}>0$. Then at $X(x_{0})$,
(2.1) $\lambda_{n}\leq\sqrt{C(n)F},\quad\lambda_{n-1}\geq\sqrt{\frac{F}{C(n)}},$
where $C(n)$ is a positive constant depending only on $n$. Furthermore, we also have
(2.2) $\dfrac{1}{C(n)F}\leq\dfrac{1}{\lambda_{n-1}\lambda_{n}}\leq\dfrac{1}{F}$
Proof.
The proof of inequality (2.1) can also be found in [13].
We just show inequality (2.2) here. First, recall the identity for $1\leq l<k\leq n$,
(2.3) $\frac{\sigma_{n-k}}{\sigma_{n-l}}(\lambda)=\dfrac{\sigma_{k}}{\sigma_{l}}(\dfrac{1}{\lambda})$
where $\lambda^{-1}=(\lambda^{-1}_{1},\lambda^{-1}_{2},\cdots,\lambda^{-1}_{n})$.
Hence, we have
$\frac{1}{F}=\sigma_{2}(\lambda^{-1})=\sum_{i=1}^{n}\sum_{k\neq i}\dfrac{1}{\lambda_{i}\lambda_{k}}.$
In addition, we have the dominant term of $\frac{1}{F}=\sigma_{2}(\lambda^{-1})$ is $\frac{1}{\lambda_{n-1}\lambda_{n}}$, thus,
$\dfrac{1}{\lambda_{n-1}\lambda_{n}}\leq\sum_{i=1}^{n}\sum_{k\neq i}\dfrac{1}{\lambda_{i}\lambda_{k}}\leq\dfrac{C(n)}{\lambda_{n-1}\lambda_{n}}.$
This completes the proof of inequality (2.2). ∎
Lemma 2.4.
Under the assumptions of Lemma 2.3, the following inequalities hold at $X(x_{0})$:
(2.4) $F^{ii}\geq\dfrac{F^{2}}{\lambda^{2}_{i}\lambda_{n}}(\text{for }i\neq n);$ $\quad F^{nn}\geq\dfrac{F^{2}}{\lambda^{2}_{i}\lambda_{n-1}}$
(2.5) $\dfrac{F}{C_{1}(n)\lambda_{n}}\leq$ $\sum_{i=1}^{n}F^{ii}\leq\dfrac{F}{\lambda_{n}}$
(2.6) $\sum_{i=1}^{n}F^{ii}h_{ii}=F$
(2.7) $\dfrac{C_{2}(n)F^{2}}{\lambda_{n}}\leq$ $\sum_{i=1}^{n}F^{ii}h^{2}_{ii}\leq\dfrac{C_{3}(n)F^{2}}{\lambda_{n}}$
where $C_{1}(n),C_{2}(n),C_{3}(n)$ are positive constants depending only on $n$.
Proof.
Equation (2.6) follows directly from formula (2.3). For inequality (2.4) and (2.5),
$F^{ii}$ $=\frac{F^{2}}{\lambda^{2}_{i}}\cdot\sum_{k\neq i}\dfrac{1}{\lambda_{k}}$
$\sum_{i=1}^{n}F^{ii}$ $=\sum_{i=1}^{n}\left(\dfrac{F^{2}}{\lambda^{2}_{i}}\cdot\sum_{k\neq i}\dfrac{1}{\lambda_{k}}\right)\leq\dfrac{F^{2}}{\lambda_{n}}\sum_{i=1}^{n}\sum_{k\neq i}\dfrac{1}{\lambda_{i}\lambda_{k}}=\dfrac{F}{\lambda_{n}},$
$\sum_{i=1}^{n}F^{ii}$ $=\sum_{i=1}^{n}\left(\dfrac{F^{2}}{\lambda^{2}_{i}}\cdot\sum_{k\neq i}\dfrac{1}{\lambda_{k}}\right)\geq\dfrac{F^{2}}{\lambda_{n}}\sum_{k\neq n}\dfrac{1}{\lambda_{n}\lambda_{k}}\geq\dfrac{F^{2}}{n\lambda_{n}}\sum_{i=1}^{n}\sum_{k\neq i}\dfrac{1}{\lambda_{i}\lambda_{k}}=\dfrac{F}{n\lambda_{n}},$
For inequality (2.7),
$\sum_{i=1}^{n}F^{ii}h^{2}_{ii}$ $=\sum_{i=1}^{n}\left(F^{2}\cdot\sum_{k\neq i}\dfrac{1}{\lambda_{k}}\right)\leq\sum_{i=1}^{n}\left(F^{2}\cdot\sum_{k\neq i}\dfrac{1}{\lambda_{n}}\right)\leq\dfrac{C_{3}(n)F}{\lambda_{n}},$
$\sum_{i=1}^{n}F^{ii}h^{2}_{ii}$ $=\sum_{i=1}^{n}\left(F^{2}\cdot\sum_{k\neq i}\dfrac{1}{\lambda_{k}}\right)\geq\sum_{i=1}^{n-1}\left(F^{2}\cdot\dfrac{1}{\lambda_{n}}\right)\geq\dfrac{C_{2}(n)F}{\lambda_{n}}.$
This completes the proof. ∎
From the formula (2.3), we see that $1/(\lambda_{n-1}\lambda_{n})$ is the dominant term in $\sigma_{2}(\lambda^{-1})=1/F$. Hence, the following bound to $\lambda_{n}$ will be useful later.
Lemma 2.5.
Under the assumptions of Lemma 2.3, at $X(x_{0})$,
$\dfrac{1}{\lambda_{1}}+\dfrac{1}{\lambda_{2}}+\cdots+\dfrac{1}{\lambda_{n-1}}\leq\frac{\lambda_{n}}{F}\leq C(n)\left(\dfrac{1}{\lambda_{1}}+\dfrac{1}{\lambda_{2}}+\cdots+\dfrac{1}{\lambda_{n-1}}\right)$
where $C(n)$ is a positive constant depending only on $n$.
Proof.
By formula (2.3), we have:
$\dfrac{1}{F}=\sigma_{2}(\lambda^{-1})=\sum_{i=1}^{n}\sum_{k\neq i}\dfrac{1}{\lambda_{i}\lambda_{k}}.$
The first inequality is immediate. For the second inequality:
$\dfrac{1}{F}=\sigma_{2}(\lambda^{-1})=\sum_{i=1}^{n-1}\sum_{k\neq i}\dfrac{1}{\lambda_{i}\lambda_{k}}+\sum_{k\neq n}\dfrac{1}{\lambda_{n}\lambda_{k}}\leq\dfrac{C_{1}(n)}{\lambda_{n-1}\lambda_{n}}+\dfrac{1}{\lambda_{n}}\sum_{k\neq n}\dfrac{1}{\lambda_{k}}\leq\dfrac{C_{2}(n)}{\lambda_{n}}\sum_{k\neq n}\dfrac{1}{\lambda_{k}}.$
where $C_{1}(n),C_{2}(n)$ are positive constants depending only on $n$. ∎
To finish this section, we demonstrate an essential concavity inequality for $F$. This inequality was introduced in [7] first and a proof can be found in [13].
Lemma 2.6.
Let $n\geq 3$ and let $F=\frac{\sigma_{n}}{\sigma_{n-2}}$ with $\lambda_{1}\geq\cdots\geq\lambda_{n}>0$. Then for any $\xi\in\mathbb{R}^{n}$, we have
(2.8) $-\sum_{i,j}F^{ii,jj}\xi_{i}\xi_{j}-\frac{F^{11}\xi_{1}^{2}}{\lambda_{1}}\geq-\frac{2}{F}\left(\sum_{i}F^{ii}\xi_{i}\right)^{2}+\frac{1}{2(n-1)}\frac{F^{11}\xi_{1}^{2}}{\lambda_{1}}.$
where $\xi_{i}=\frac{\partial\xi}{\partial x_{i}}$.
3. Jacobi inequality
To establish curvature estimates, the primary challenges arise from the first and second derivatives of curvature. In this section, we generalize the Jacobi inequality in [13] to Riemannian manifolds using the concavity inequality in Lemma 2.6 in order to eliminate the second derivatives of curvature subsequently.
For simplicity, assume $\lambda_{1}$ is sufficiently large. Let $M\subset\mathbb{R}^{n+1}$ be a hypersurface represented as a graph $X=(x,u(x))$ over $B_{r}\subseteq\mathbb{R}^{n}$ satisfying the equation (1.1). At any point $x\in B_{r}$, let $h=(h_{ij})$ denote the second fundamental form of the graph and $\lambda=(\lambda_{1},\lambda_{2},\cdots,\lambda_{n})$ be its principal curvatures.
We now recall several fundamental identities in Riemannian geometry. First, we choose an orthonormal frame in $\mathbb{R}^{n+1}$ such that $\{e_{1},e_{2},\cdots,e_{n}\}$ are tangent to $M$ and $\nu$ is the outer normal to $M$. The following fundamental identities hold:
[…]
where $R_{ijkl}$ is the curvature tensor. Define $h_{ijkl}=\nabla_{l}\nabla_{k}h_{ij}$, the commutator formula yields:
(3.5) $h_{ijkl}-h_{ijlk}=\sum_{m}h_{im}R_{ijmk}+\sum_{m}h_{mj}R_{mikl}.$
Combine equation (3.5) with Codazzi and Gauss equation, we derive:
(3.6) $h_{iikk}$ $=h_{kkii}+\sum_{m}(h_{im}h_{mi}h_{kk}-h^{2}_{mk}h_{ii}).$
We now proceed to generalize the Jacobi inequality.
Lemma 3.1.
For $n\geq 3$, suppose $f\in C^{1,1}(B_{r})$ be a positive function and let $M$ be a convex graph over $B_{r}\subset\mathbb{R}^{n}$ solving equation (1.1). At any $X(x_{0}),x_{0}\in B_{r}$, assume $h$ is diagonalized with ordered principal curvatures $\lambda_{1}\geq\cdots\geq\lambda_{n}$. Define $b=\ln\lambda_{1}$. Then, at $X(x_{0})$, the following inequality holds in the viscosity sense:
(3.7) $\sum_{i}F^{ii}b_{ii}\geq c(n)\sum_{i}F^{ii}b_{i}^{2}+\sum_{i}F^{ii}h_{ii}h_{11}-\sum_{i}F^{ii}h^{2}_{ii}-C,$
where $c(n)$ is a positive constant depending only on $n$, and $C$ depends only on $n$, $\|M\|_{C^{0,1}\left(B_{r}\right)}$, $\min_{B_{r}}f$, and $\|f\|_{C^{1,1}\left(B_{r}\right)}$.
Proof.
Suppose $\lambda_{1}$ has multiplicity $m$ at $X(x_{0})$. By Lemma 5 in [1], we have
(3.8) $\delta_{\alpha\beta}\cdot{(\lambda_{1})}_{i}=h_{\alpha\beta i},\quad 1\leq\alpha,\beta\leq m,$
(3.9) ${(\lambda_{1})}_{ii}\geq h_{11ii}+2\sum_{p>m}\frac{h_{1pi}^{2}}{\lambda_{1}-\lambda_{p}},$
in the viscosity sense.
It follows that
$b_{ii}=\frac{{(\lambda_{1})}_{ii}}{\lambda_{1}}-\frac{{(\lambda_{1})}_{i}^{2}}{\lambda_{1}^{2}}\geq\frac{h_{11ii}}{\lambda_{1}}+2\sum_{p>m}\frac{h_{1pi}^{2}}{\lambda_{1}(\lambda_{1}-\lambda_{p})}-\frac{h_{11i}^{2}}{\lambda_{1}^{2}}.$
Contracting with $F^{ii}$,
(3.10) $\sum_{i}F^{ii}b_{ii}\geq\sum_{i}\frac{F^{ii}h_{11ii}}{\lambda_{1}}+2\sum_{i}\sum_{p>m}\frac{F^{ii}h_{1pi}^{2}}{\lambda_{1}(\lambda_{1}-\lambda_{p})}-\sum_{i}\frac{F^{ii}h_{11i}^{2}}{\lambda_{1}^{2}}.$
Differentiate equation (1.1) twice with respect to $e_{1}$:
(3.11) $F^{ii}(\lambda_{i})_{1}$ $=d_{X}f(e_{1}),$
(3.12) $F^{ii}(\lambda_{i})_{11}+F^{pq,rs}h_{pq1}h_{rs1}$ $=d_{X,X}f(e_{1},e_{1})\geq-C\lambda_{1}-C,$
where $C$ is a positive constant depending only on $n$, $\|M\|_{C^{0,1}\left(B_{r}\right)}$, $\min_{B_{r}}f$, and $\|f\|_{C^{1,1}\left(B_{r}\right)}$ and may change from line to line.
Substituting (3.6) and (3.12) into (3.10),
$\sum_{i}F^{ii}b_{ii}\geq$ $-\sum_{p,q,r,s}\frac{F^{pq,rs}h_{pq1}h_{rs1}}{\lambda_{1}}+2\sum_{i}\sum_{p>m}\frac{F^{ii}h_{1pi}^{2}}{\lambda_{1}(\lambda_{1}-\lambda_{p})}-\sum_{i}\frac{F^{ii}h_{11i}^{2}}{\lambda_{1}^{2}}$
(3.13) $+\sum_{i}F^{ii}h_{ii}h_{11}-\sum_{i}F^{ii}h^{2}_{ii}-C.$
By Lemma 2.2,
$-\sum_{p,q,r,s}F^{pq,rs}u_{pq1}u_{rs1}=-\sum_{i,j}F^{ii,jj}u_{ii1}u_{jj1}-\sum_{i\neq j}F^{ij,ji}u_{ij1}^{2}.$
Substituting into (3),
(3.14) $\sum_{i}F^{ii}b_{ii}\geq$ $\ -\sum_{i,j}\frac{F^{ii,jj}h_{ii1}h_{jj1}}{\lambda_{1}}-\sum_{i\neq j}\frac{F^{ij,ji}h_{ij1}^{2}}{\lambda_{1}}+2\sum_{i}\sum_{p>m}\frac{F^{ii}h_{1pi}^{2}}{\lambda_{1}(\lambda_{1}-\lambda_{p})}$
$\ -\sum_{i}\frac{F^{ii}h_{11i}^{2}}{\lambda_{1}^{2}}+\sum_{i}F^{ii}h_{ii}h_{11}-\sum_{i}F^{ii}h^{2}_{ii}-C.$
By Lemma 2.2,
$-\sum_{i\neq j}\frac{F^{ij,ji}h_{ij1}^{2}}{\lambda_{1}}\geq-2\sum_{i>m}\frac{F^{i1,1i}h_{i11}^{2}}{\lambda_{1}}=2\sum_{i>m}\frac{(F^{ii}-F^{11})h_{11i}^{2}}{\lambda_{1}(\lambda_{1}-\lambda_{i})}.$
On the other hand,
$2\sum_{i}\sum_{p>m}\frac{F^{ii}h_{1pi}^{2}}{\lambda_{1}(\lambda_{1}-\lambda_{p})}\geq 2\sum_{p>m}\frac{F^{11}h_{1p1}^{2}}{\lambda_{1}(\lambda_{1}-\lambda_{p})}=2\sum_{i>m}\frac{F^{11}h_{11i}^{2}}{\lambda_{1}(\lambda_{1}-\lambda_{i})}.$
Combining the above two inequalities and substituting into inequality (3.14),
$\sum_{i}F^{ii}b_{ii}\geq$ $\ -\sum_{i,j}\frac{F^{ii,jj}h_{ii1}h_{jj1}}{\lambda_{1}}+2\sum_{i>m}\frac{F^{ii}h_{11i}^{2}}{\lambda_{1}(\lambda_{1}-\lambda_{i})}-\sum_{i}\frac{F^{ii}h_{11i}^{2}}{\lambda_{1}^{2}}$
(3.15) $+\sum_{i}F^{ii}h_{ii}h_{11}-\sum_{i}F^{ii}h^{2}_{ii}-C.$
By equation (3.8) and Codazzi equation,
(3.16) $h_{11i}=h_{1i1}=\delta_{1i}\cdot(\lambda_{1})_{1}=0,\quad 1<i\leq m.$
Substituting into inequality (3),
$\sum_{i}F^{ii}b_{ii}\geq$ $-\sum_{i,j}\frac{F^{ii,jj}h_{ii1}h_{jj1}}{\lambda_{1}}+\sum_{i>m}\frac{F^{ii}h_{11i}^{2}}{\lambda_{1}^{2}}-\frac{F^{11}h_{111}^{2}}{\lambda_{1}^{2}}$
$+\sum_{i}F^{ii}h_{ii}h_{11}-\sum_{i}F^{ii}h^{2}_{ii}-C.$
Together with Lemma 2.6 and equation (3.16), then:
$\sum_{i}F^{ii}b_{ii}\geq$ $-\frac{2}{\lambda_{1}F}\left(\sum_{i}F^{ii}h_{ii1}\right)^{2}+\sum_{i>m}\frac{F^{ii}h_{11i}^{2}}{\lambda_{1}^{2}}+c(n)\frac{F^{11}h_{111}^{2}}{\lambda_{1}^{2}}$
$+\sum_{i}F^{ii}h_{ii}h_{11}-\sum_{i}F^{ii}h^{2}_{ii}-C$
$\geq$ $\ -\frac{2f_{1}^{2}}{\lambda_{1}F}+c(n)\sum_{i}\frac{F^{ii}h_{11i}^{2}}{\lambda_{1}^{2}}+\sum_{i}F^{ii}h_{ii}h_{11}-\sum_{i}F^{ii}h^{2}_{ii}-C$
$\geq$ $\ c(n)\sum_{i}F^{ii}b_{i}^{2}+\sum_{i}F^{ii}h_{ii}h_{11}-\sum_{i}F^{ii}h^{2}_{ii}-C.$
The lemma is now proved.
4. Proof of the Theorem 1.1
In this section, we complete the proof of Theorem 1.1 using an appropriate auxiliary function and subsequent calculations. Additionally, we employ the Einstein summation convention to simplify the notation, unless explicitly stated otherwise.
Proof.
Without loss of generality, we assume $r=1$. The argument can be extended naturally to general $r>0$ through rescaling. To maintain consistency, we use the same assumption as Section 3. Then, we use the same auxiliary function in $B_{1}$ as Guan, Qiu’s [6], that is:
(4.1) $P(X(x))=2\log\rho(X)+\log\log\lambda_{1}-\beta\frac{(X,\nu)}{(\nu,E_{n+1})}+\alpha\frac{1}{(\nu,E_{n+1})^{2}},$
where $E_{n+1}=(0,\cdots,0,1)$, $\rho(X)=1-|X|_{\mathbb{R}^{n+1}}^{2}+(X,E_{n+1})^{2}=1-|x|_{\mathbb{R}^{n}}^{2}$, $\alpha,\beta$ are positive constants to be determined. Consequently, the maximum of $P$ is attained at an interior point $x_{0}$ of $B_{1}$.
Since $|(\nu,E_{n+1})|$ is bounded above and below for $x\in B_{1}$ under the assumption that $M$ is a graph and $\lambda_{1}$ is sufficiently large, we obtain:
$C$ $\leq\rho^{2}(x)\cdot\log\lambda_{1},$
that is,
(4.2) $\rho^{2}$ $\geq\frac{C}{\log\lambda_{1}},\quad\rho\geq\frac{C}{\sqrt{\log\lambda_{1}}}.$
where $C$ is a positive constant dependent only on $n$, $|M|_{C^{1}\left(B_{1}\right)}$, $\min_{B_{1}}|f|$, and $\|f\|_{C^{2}\left(B_{1}\right)}$, which may change between steps.
Differentiating $P(X)$ at $x_{0}\in B_{1}$, we obtain the first derivative:
(4.3) 00 0 $=P_{i}=\frac{2\rho_{i}}{\rho}+\frac{h_{11i}}{h_{11}\log h_{11}}-\beta\left[\frac{(X,\nu)}{(\nu,E_{n+1})}\right]_{i}-\alpha\frac{2(\nu_{i},E_{n+1})}{(\nu,E_{n+1})^{3}}.$
Since $h$ is diagonalized at $x_{0}$, the second derivative satisfies:
$0\geq F^{ij}P_{ij}=$ $\frac{2F^{ii}\rho_{ii}}{\rho}-\frac{2F^{ii}\rho_{i}^{2}}{\rho^{2}}-\beta F^{ii}\left[\frac{(X,\nu)}{(\nu,E_{n+1})}\right]_{ii}+\frac{F^{ii}(\lambda_{1})_{ii}}{h_{11}\log h_{11}}$
(4.4) $-\left(\log\lambda_{1}+1\right)\frac{F^{ii}h^{2}_{11i}}{h^{2}_{11}\log h^{2}_{11}}-\alpha\frac{2F^{ii}(\nu_{ii},E_{n+1})}{(\nu,E_{n+1})^{3}}+6\alpha\frac{F^{ii}(\nu_{i},E_{n+1})^{2}}{(\nu,E_{n+1})^{4}}.$
As noted in Section 3, the critical barrier in the estimate lies in controlling the derivative of curvature. To manage these terms we express certain elements in inequality (4.4) using derivatives of curvature:
(4.5) $\nu_{ii}$ $=\left(\sum_{k}h_{ik}e_{k}\right)_{i}=\sum_{k}h_{iik}e_{k}-\sum_{k}h_{ik}h_{ki}\nu,$
(4.6) $\left[\frac{(X,\nu)}{(\nu,E_{n+1})}\right]_{i}$ $=\sum_{l}\frac{(X,e_{l})h_{il}}{(\nu,E_{n+1})}-\sum_{l}\frac{(e_{l},E_{n+1})h_{il}(X,\nu)}{(\nu,E_{n+1})^{2}},$
$\left[\frac{(X,\nu)}{(\nu,E_{n+1})}\right]_{ii}$ $=\frac{h_{ii}}{(\nu,E_{n+1})}-2\left[\frac{(X,\nu)}{(\nu,E_{n+1})}\right]_{i}\cdot\sum_{l}\frac{(e_{l},E_{n+1})h_{il}}{(\nu,E_{n+1})^{2}}$
(4.7) $\quad+\sum_{l}\frac{(X,e_{l})(\nu,E_{n+1})-(e_{l},E_{n+1})(X,\nu)}{(\nu,E_{n+1})^{2}}h_{iil},$
(4.8) $\rho_{i}$ $=-2(X,e_{i})+2(X,E_{n+1})(e_{i},E_{n+1}),$
(4.9) $\rho_{ii}$ $=-2+2(X,\nu)h_{ii}+2(e_{i},E_{n+1})^{2}-2(X,E_{n+1})h_{ii}(\nu,E_{n+1}),$
Insert equation (4.54.9 and 4.3) into (4.4), we obtain:
$F^{ij}P_{ij}\geq$ $-\dfrac{4\sum_{i}F^{ii}}{\rho}+\frac{4F^{ii}h_{ii}[(X,\nu)-(X,E_{n+1})(\nu,E_{n+1})]}{\rho}$
$+\frac{F^{ii}(\lambda_{1})_{ii}}{h_{11}\log h_{11}}-\frac{8F^{ii}[(X,e_{i})-(X,E_{n+1})(e_{i},E_{n+1})]^{2}}{\rho^{2}}$
$-(\log\lambda_{1}+1)\frac{F^{ii}h^{2}_{iij}}{h_{11}^{2}\log^{2}h_{11}}+2F^{ii}\dfrac{(e_{i},E_{n+1})h_{ii}}{(\nu,E_{n+1})}\left(\dfrac{2\rho_{i}}{\rho}+\frac{h_{11i}}{h_{11}\log h_{11}}\right)$
$-\frac{\alpha\ F^{ii}h_{iik}(e_{i},E_{n+1})}{(\nu,E_{n+1})^{3}}-\beta\dfrac{F^{ii}h_{ii}}{(\nu,E_{n+1})}$
(4.10) $-\beta F^{ii}h_{iil}\frac{(X,e_{l})(e_{i},E_{n+1})-(e_{l},E_{n+1})(X,\nu)}{(\nu,E_{n+1})^{2}}+\dfrac{2\alpha F^{ij}h^{2}_{ii}}{(\nu,E_{n+1})^{2}},$
Next, we analyze some terms in inequality (4.10). By Equation (2.6), we have:
$\frac{4F^{ii}h_{ii}[(X,\nu)-(X,E_{n+1})(\nu,E_{n+1})]}{\rho}=$ $-\dfrac{4F[(X,\nu)-(X,E_{n+1})(\nu,E_{n+1})]}{\rho}$
(4.11) $\geq$ $-\dfrac{C(\parallel f\parallel_{C^{2}},|M|_{C_{1}})}{\rho}.$
Recalling that we define $b=\ln\lambda_{1}$, we apply the Jacobi inequality to eliminate the second derivative term of curvature, we obtain:
$\dfrac{F^{ii}(\lambda_{1})_{ii}}{h_{11}\log h_{11}}$ $=\frac{F^{ii}b_{ii}}{\log h_{11}}+\frac{1}{\log h_{11}}\cdot\frac{F^{ii}(\lambda_{1})_{i}^{2}}{\lambda_{1}^{2}}$
$\geq\frac{1}{\log h_{11}}\left(C(n)F^{ii}b_{i}^{2}+Fh_{11}-\sum_{i}F^{ii}h_{ii}^{2}-C\right)$
(4.12) $=\frac{C(n)}{\log h_{11}}\cdot\frac{F^{ii}(\lambda_{1})_{i}^{2}}{\lambda_{1}^{2}}+\frac{Fh_{11}}{\log h_{11}}-\frac{\sum_{i}F^{ii}h_{ii}^{2}}{\log h_{11}}-C.$
Combine inequality (4.104.12) and introducing $a_{i}=[(X,e_{i})-(X,E_{n+1})(e_{i},E_{n+1})]$, we obtain:
$F^{ij}P_{ij}$ $\geq-\dfrac{4\sum_{i}F^{ii}}{\rho}-\frac{C}{\rho}-C-\frac{8\sum_{i}F^{ii}\left[(X,e_{i})-(X,E_{n+1})(e_{i},E_{n+1})\right]^{2}}{\rho^{2}}$
$\quad+\frac{C(n)}{\log h_{11}}\cdot\frac{F^{ii}(\lambda_{1})_{i}^{2}}{\lambda_{1}^{2}}+\frac{Fh_{11}}{\log h_{11}}-\frac{\sum_{i}F^{ii}h_{ii}^{2}}{\log h_{11}}-(\log h_{ii}+1)\dfrac{F^{ii}h^{2}_{11i}}{h^{2}_{11}\log^{2}h_{11}}$
$\quad+\frac{2F^{ii}h_{ii}(e_{i},E_{n+1})}{(\nu,E_{n+1})^{2}}\left(\frac{2\rho_{i}}{\rho}+\frac{h_{11i}}{h_{11}\log h_{11}}\right)+2\alpha\frac{F^{ii}h_{ii}^{2}}{(\nu,E_{n+1})^{2}}-C\beta-C\alpha.$
$\geq-\dfrac{4\sum_{i}F^{ii}}{\rho}-C(\frac{1}{\rho}+\beta+\alpha)-\dfrac{8F^{ii}a_{i}^{2}}{\rho^{2}}+\frac{C(n)}{\log h_{11}}\cdot\frac{F^{ii}(\lambda_{1})_{i}^{2}}{\lambda_{1}^{2}}+\frac{Fh_{11}}{\log h_{11}}$
$\quad-\frac{F^{ii}h_{ii}^{2}}{\log h_{11}}-(\log h_{11}+1)\dfrac{F^{ii}h^{2}_{11i}}{h^{2}_{11}\log^{2}h_{11}}+\frac{2F^{ii}h_{ii}(e_{i},E_{n+1})}{(\nu,E_{n+1})^{2}}\left(\frac{2\rho_{i}}{\rho}+\frac{h_{11i}}{h_{11}\log h_{11}}\right)$
(4.13) $\quad+2\alpha\frac{F^{ii}h_{ii}^{2}}{(\nu,E_{n+1})^{2}}.$
By Cauchy-Schwarz inequality, we have
(4.14) $4F^{ii}\frac{h_{ii}\rho_{i}}{\rho(\nu,E_{n+1})}$ $\geq-\frac{2F^{ii}h_{ii}^{2}}{(\nu,E_{n+1})^{2}}-\frac{2F^{ii}\rho_{i}^{2}}{\rho^{2}},$
(4.15) $2F^{ii}\frac{h_{11i}h_{ii}}{(\nu,E_{n+1})h_{11}\log h_{11}}$ $\geq-\frac{F^{ii}h_{ii}^{2}}{(\nu,E_{n+1})^{2}}-\frac{F^{ii}h_{11i}^{2}}{h_{11}^{2}\log^{2}h_{11}}.$
Thus, substitute inequality (4.14) and (4.15) into (4.13), we attain
$F^{ij}P_{ij}$ $\geq-\dfrac{4\sum_{i}F^{ii}}{\rho}-\dfrac{16F^{ii}a_{i}^{2}}{\rho^{2}}-C(\frac{1}{\rho}+\beta+\alpha)+\frac{C(n)}{\log h_{11}}\cdot\frac{F^{ii}(\lambda_{1})_{i}^{2}}{\lambda_{1}^{2}}+\frac{Fh_{11}}{\log h_{11}}$
(4.16) $\quad-\frac{F^{ii}h_{ii}^{2}}{\log h_{11}}-(\log h_{11}+2)\dfrac{F^{ii}h^{2}_{11i}}{h^{2}_{11}\log^{2}h_{11}}+(2\alpha-3)\frac{F^{ii}h_{ii}^{2}}{(\nu,E_{n+1})^{2}}.$
Since we can choose $h_{11},\alpha$ to be large enough, and applying formula (3.9), we obtain:
$0\geq F^{ij}P_{ij}$ $\geq-\dfrac{4\sum_{i}F^{ii}}{\rho}-\dfrac{16F^{ii}a_{i}^{2}}{\rho^{2}}-C(\frac{1}{\rho}+\beta+\alpha)+\frac{F^{ii}h_{11i}^{2}}{20h_{11}^{2}\log h_{11}}$
(4.17) $\quad+\frac{Fh_{11}}{\log h_{11}}+\alpha\frac{F^{ii}h_{ii}^{2}}{(\nu,E_{n+1})^{2}}.$
Thus, we arrive at the required inequality, as there are no negative terms of derivative terms of curvature in this inequality. Examining inequality (4.17), the dominant negative terms are the first and second terms. The consequent discussion on $a_{i}$ and $\rho$ provides a range of these two negative terms which will be useful later.
Thus far, all computations have been performed using a local frame on the hypersurface. We now switch to an orthonormal coordinate system $\{E_{1},E_{2},\cdots,E_{n},E_{n+1}\}$ in $\mathbb{R}^{n+1}$, where $E_{i}\perp E_{n+1}$ for all $i\in\{1,\cdots,n\}$. In this new coordinate, we can decompose the vector $X$ as follows:
(4.18) $X=\sum_{i=1}^{n}(X,E_{i})E_{i}+(X,E_{n+1})E_{n+1}.$
Thus,
(4.19) $\rho=1-\sum_{i=1}^{n}(X,E_{i})^{2}.$
Recalling that
(4.20) $a_{j}=(X,e_{j})-(X,E_{n+1})(e_{j},E_{n+1})=\sum_{i}(X,E_{i})(E_{i},e_{j}).$
We obtain:
$\sum_{i=1}^{n}a_{i}^{2}$ $=$ $\sum_{i,k,l=1}^{n}(X,E_{k})(X,E_{l})(E_{k},e_{i})(E_{l},e_{i})$
$=$ $\sum_{k,l=1}^{n}(X,E_{k})(X,E_{l})(\delta_{kl}-(E_{k},\nu)(E_{l},\nu)).$
Since $(\nu,E_{n+1})$ is bounded from above and below, we obtain:
(4.21) $\frac{1}{C}\sum_{i=1}^{n}(X,E_{i})^{2}\leq\sum_{i=1}^{n}a_{i}^{2}\leq C\sum_{i=1}^{n}(X,E_{i})^{2}\leq C(|M|_{C_{1}},n).$
Combine relation (4.19), (4.21), and (4.2), we derive:
(4.22) $\dfrac{1}{C}(1-\rho)\leq\sum_{i=1}^{n}a^{2}_{i}\leq C(1-\rho)\leq C(1-\dfrac{1}{\sqrt{\log\lambda_{1}}})\leq C(|M|_{C_{1}},n).$
Therefore, our proof strategy is to use positive terms in inequality (4.17) to control those two dominant negative terms. We categorize the proof into three cases based on the values of $\rho$ and $a_{i}$.
Case 1: $\sum_{i=1}^{n}|(X,E_{i})|^{2}\leq\frac{1}{2}$
In this case, $\rho\geq\frac{1}{2}$. Combine inequality (4.17) with (2.5), (2.7), and (4.22), we obtain:
(4.23) $F^{ij}P_{ij}$ $\geq-\frac{C}{\lambda_{n}}-C\left(1+\beta+\alpha\right)+\frac{Fh_{11}}{\log h_{11}}+\frac{C\alpha}{\lambda_{n}}.$
Choosing $\alpha$ sufficiently large, we obtain:
(4.24) $F^{ij}P_{ij}$ $\geq-C\left(1+\beta+\alpha\right)+\frac{Fh_{11}}{\log h_{11}}.$
This provides the estimate for this case. Hence, it remains to discuss $\sum_{i=1}^{n}|(X,E_{i})|^{2}>\frac{1}{2}$.
Case 2: $\sum_{i=1}^{n}|(X,E_{i})|^{2}>\frac{1}{2}$ and suppose $\sqrt{\log\lambda_{1}}\geq C\lambda_{n-1}$
By lemma 2.5, we have:
$\left(\frac{1}{\lambda_{n-1}}+\dots+\frac{1}{\lambda_{1}}\right)\leq\dfrac{\lambda_{n}}{F}\leq C(n)\left(\frac{1}{\lambda_{n-1}}+\dots+\frac{1}{\lambda_{1}}\right)$
Using the condition of case 2, we obtain:
(4.25) $\dfrac{\lambda_{n}}{F}\geq\frac{1}{\lambda_{n-1}}+\dots+\frac{1}{\lambda_{1}}\geq\dfrac{1}{\lambda_{n-1}}\geq\frac{C}{\sqrt{\log h_{11}}},$
Substituting relation (4.2) and (4.25) into (4.17), we derive:
$F^{ij}P_{ij}\geq$ $-\frac{4\sum_{i}F^{ii}}{\rho}-\frac{16F^{ii}a_{i}^{2}}{\rho^{2}}-C\left(\frac{1}{\rho}+\beta+\alpha\right)+\frac{F^{ii}h_{11i}^{2}}{20h^{2}_{11}\log h_{11}}$
$+\dfrac{Fh_{11}}{\log h_{11}}+\alpha\dfrac{F^{ii}h^{2}_{ii}}{(\nu,E_{n+1})^{2}}$
$\geq$ $-\frac{C}{\lambda_{n}\rho}-\frac{C}{\lambda_{n}\rho^{2}}-C\left(\frac{1}{\rho}+\beta+\alpha\right)+\frac{Fh_{11}}{\log h_{11}}$
(4.26) $\geq$ $-C\log\lambda_{1}\cdot\sqrt{\log\lambda_{1}}-C\left(1+\beta+\alpha\right)+\frac{Fh_{11}}{\log h_{11}}$
For sufficiently large $\lambda_{1}$, choosing $\frac{1}{2}F\lambda_{1}>C(\log\lambda_{1})^{\frac{5}{2}}$, gives the estimate for case 2.
Case 3.1: If $\sqrt{\log\lambda_{1}}<C\lambda_{n-1}$, and $\sum_{i=1}^{n}|(X,E_{i})|^{2}>\frac{1}{2}$. Denote $d=\sqrt{\frac{1}{2Cn}}$ and assume $|a_{n}|<d$, where the $C$ here is the one in relation (4.21).
Given this condition, there must exist some $1\leq j<n$ such that $|a_{j}|>d$. Selecting such an index $j$, we analyze the following term from inequality (4.17):
(4.27) $\frac{F^{ii}h_{11i}^{2}}{20h^{2}_{11}\log h_{11}}=\sum_{i\neq j}\frac{F^{ii}h_{11i}^{2}}{20h^{2}_{11}\log h_{11}}+\frac{F^{jj}h_{11j}^{2}}{20h^{2}_{11}\log h_{11}}\geq\frac{F^{jj}h_{11j}^{2}}{20h^{2}_{11}\log h_{11}}$
Here, $j$ is fixed throughout this expression, so $F^{jj}h^{2}_{11j}$ does not represent a summation.
Applying triangle inequality to Equation (4.3), then at the critical point $x_{0}$,
$\left|\frac{h_{11j}}{h_{11}\log h_{11}}\right|=$ $\left|\frac{4a_{j}}{\rho}+\beta\frac{b_{j}h_{jj}}{(\nu,E_{n+1})}+\alpha\frac{2h_{jj}(e_{j},E_{n+1})}{(\nu,E_{n+1})^{3}}\right|$
(4.28) $\geq$ $\left|\left|\frac{4a_{j}}{\rho}\right|-\left|\frac{\beta b_{j}}{(\nu,E_{n+1})}+\frac{2\alpha(e_{j},E_{n+1})}{(\nu,E_{n+1})^{3}}\right|h_{jj}\right|$
where $b_{j}=(X,e_{j})(\nu,E_{n+1})-(E_{n+1},e_{j})(X,\nu)$.
Substitute this condition along with inequality (4) into (4.27) and using inequality (2.4), we obtain:
$\frac{F^{jj}h_{11j}^{2}}{20h^{2}_{11}\log h_{11}}\geq$ $\frac{CF^{2}\log h_{11}}{\lambda_{j}^{2}\lambda_{n}}\cdot\left(\frac{Cd^{2}}{\rho^{2}}+\left|\frac{\beta b_{j}}{(\nu,E_{n+1})}+\frac{2\alpha(e_{j},E_{n+1})}{(\nu,E_{n+1})^{3}}\right|^{2}h_{jj}^{2}\right.$
$\hskip 85.35826pt-\left.Cd\cdot{\left|\frac{\beta b_{j}}{(\nu,E_{n+1})}+\frac{2\alpha(e_{j},E_{n+1})}{(\nu,E_{n+1})^{3}}\right|\frac{h_{jj}}{\rho}}\right)$
$\geq$ $\frac{Cd^{2}F^{2}\log h_{11}}{\lambda_{j}^{2}\lambda_{n}\rho^{2}}+\frac{CF^{2}\log h_{11}}{\lambda_{n}}\cdot\left|{\frac{\beta b_{j}}{(\nu,E_{n+1})}}+{\frac{2\alpha(e_{j},E_{n+1})}{(\nu,E_{n+1})^{3}}}\right|^{2}$
$-\frac{CdF^{2}\log h_{11}}{\lambda_{n}}\cdot\left|\frac{\beta b_{j}}{(\nu,E_{n+1})}+\frac{2\alpha(e_{j},E_{n+1})}{(\nu,E_{n+1})^{3}}\right|$
(4.29) $\geq$ $\left(\left|\frac{\beta b_{j}}{(\nu,E_{n+1})}+\frac{2\alpha(e_{j},E_{n+1})}{(\nu,E_{n+1})^{3}}\right|^{2}-d\left|\frac{\beta b_{j}}{(\nu,E_{n+1})}+\frac{2\alpha(e_{j},E_{n+1})}{(\nu,E_{n+1})^{3}}\right|\right)\frac{CF^{2}\log h_{11}}{\lambda_{n}}.$
Assume first that $\alpha$ is sufficiently large so that Case 1 applies. We then analyze the coefficient on the right-hand side of the inequality:
In fact, we can decompose vector $\nu$ as following
(4.30) $\nu=\sum_{i}(E_{i},\nu)E_{i}+(E_{n+1},\nu)E_{n+1}.$
and for any $l\in\{1,\cdots,n\}$, we have:
$b_{l}$ $=\sum_{i}[(X,E_{i})(E_{i},e_{l})(\nu,E_{n+1})-(E_{n+1},e_{l})(X,E_{i})(E_{i},\nu)]$
$\quad+(E_{n+1},e_{l})[(X,E_{n+1})(\nu,E_{n+1})-(X,E_{n+1})(E_{n+1},\nu)]$
(4.31) $=a_{l}(\nu,E_{n+1})-(E_{n+1},e_{l})\sum_{i=1}^{n}(X,E_{i})(E_{i},\nu).$
① If $(e_{j},E_{n+1})=0$, then from Equation (4.31), we obtain:
(4.32) $|b_{j}|=|a_{j}(\nu,E_{n+1})|\geq d\cdot|(\nu,E_{n+1})|.,$
Consequently,
(4.33) $\quad\left|\frac{\beta b_{j}}{(\nu,E_{n+1})}+\frac{2\alpha(e_{j},E_{n+1})}{(\nu,E_{n+1})^{3}}\right|=\left|\frac{\beta b_{j}}{(\nu,E_{n+1})}\right|\geq\beta d.$
Then we choose $\beta\geq 2$ to obtain:
(4.34) $\quad\left|\frac{\beta b_{j}}{(\nu,E_{n+1})}\right|^{2}-d\left|\frac{\beta b_{j}}{(\nu,E_{n+1})}\right|\geq\frac{1}{2}\beta^{2}d^{2}=B_{1},$
② If $|(e_{j},E_{n+1})|\geq A>0$, where $A\in(0,1]$ is a positive constant, then since $|b_{j}|<C(|M|_{C^{1}(B_{1})},n)$, by firstly choosing $\beta\geq 2$ and then $\alpha\geq\frac{2\beta(\|M\|_{C^{1}(B_{r})}+d)}{A}$ we have:
(4.35) $\quad\left|\frac{\beta b_{j}}{(\nu,E_{n+1})}+\frac{2\alpha(e_{j},E_{n+1})}{(\nu,E_{n+1})^{3}}\right|\geq\left|\frac{\alpha A}{(\nu,E_{n+1})^{3}}\right|$
Furthermore, we choose
$\alpha\geq max\{\frac{2\beta(\|M\|_{C^{1}(B_{r})}+d)}{A},\frac{8d}{A^{2}},\frac{2\sqrt{d\beta(\|M\|_{C^{1}(B_{r})}+d)}}{A}\}$
then to ensure:
(4.36) $\quad\left|\frac{\beta b_{j}}{(\nu,E_{n+1})}+\frac{2\alpha(e_{j},E_{n+1})}{(\nu,E_{n+1})^{3}}\right|^{2}-d\left|\frac{\beta b_{j}}{(\nu,E_{n+1})}+\frac{2\alpha(e_{j},E_{n+1})}{(\nu,E_{n+1})^{3}}\right|\geq B_{2},$
where
$B_{2}=\frac{1}{2}\left(\frac{\alpha A}{(\nu,E_{n+1})^{3}}\right)^{2}.$
Therefore, if $A\in(0,1]$ exists, we set $B=B_{2}$, otherwise we let $B=B_{1}$. However, we can always choose $B$ large enough by choose $\alpha,\beta$ large enough in both cases.
By the analysis above, inequality (4) becomes:
(4.37) $\quad\frac{F^{jj}h_{11j}^{2}}{20h^{2}_{11}\log h_{11}}\geq\frac{CBF^{2}\log h_{11}}{\lambda_{n}}$
Plugging relation (4.37) and (4.2) into (4.17), we obtain:
$F^{ij}P_{ij}$ $\geq\frac{4\sum_{i}F^{ii}}{\rho}-\frac{16F^{ii}a_{i}^{2}}{\rho^{2}}-C\left(\frac{1}{\rho}+\beta+\alpha\right)+\frac{F^{ii}h_{11i}^{2}}{20h^{2}_{11}\log h_{11}}$
$\quad+\frac{Fh_{11}}{\log h_{11}}+\alpha\frac{F^{ii}h_{ii}^{2}}{(\nu,E_{n+1})^{2}}$
$\geq-\frac{C}{\lambda_{n}\rho}-\frac{C}{\lambda_{n}\rho^{2}}-C\left(\frac{1}{\rho}+\beta+\alpha\right)+\frac{CBF^{2}\log h_{11}}{\lambda_{n}}$
$\quad+\frac{Fh_{11}}{\log h_{11}}+\alpha\frac{F^{ii}h_{ii}^{2}}{(\nu,E_{n+1})^{2}}$
(4.38) $\geq-\frac{C\log h_{11}}{\lambda_{n}}-C\left(1+\beta+\alpha\right)+\frac{CBF^{2}\log h_{11}}{\lambda_{n}}+\frac{Fh_{11}}{\log h_{11}}$
Choose $\alpha$ and $\beta$ as stated above to ensure $B$ large enough, we have the estimate for this case.
Case 3.2: If $\sqrt{\log\lambda_{1}}<C\lambda_{n-1}$, and $\sum_{i=1}^{n}|(X,E_{i})|^{2}>\frac{1}{2}$. Assume only $|a_{n}|\geq d$, where $d$ is the same as in case 3.1.
Similarly to case 3.1, we obtain
$\frac{F^{nn}h_{11n}^{2}}{20h^{2}_{11}\log h_{11}}$ $\geq\frac{CF^{2}\log h_{11}}{\lambda_{n}^{2}\lambda_{n-1}}\cdot\left(\frac{Cd^{2}}{\rho^{2}}+\left|\frac{\beta b_{n}}{(\nu,E_{n+1})}+\frac{2\alpha(e_{n},E_{n+1})}{(\nu,E_{n+1})^{3}}\right|^{2}h_{nn}^{2}\right.$
$\hskip 105.2751pt\left.-Cd\left|\frac{\beta b_{n}}{(\nu,E_{n+1})}+\frac{2\alpha(e_{n},E_{n+1})}{(\nu,E_{n+1})^{3}}\right|\frac{h_{nn}}{\rho}\right)$
$=\frac{Cd^{2}F^{2}\log h_{11}}{\lambda_{n}^{2}\lambda_{n-1}\rho^{2}}+\frac{CF^{2}\log h_{11}}{\lambda_{n-1}}\left|\frac{\beta b_{n}}{(\nu,E_{n+1})}+\frac{2\alpha(e_{n},E_{n+1})}{(\nu,E_{n+1})^{3}}\right|^{2}$
$\quad-\frac{CdF^{2}\log h_{11}}{\lambda_{n}\lambda_{n-1}\rho}\left|\frac{\beta b_{n}}{(\nu,E_{n+1})}+\frac{2\alpha(e_{n},E_{n+1})}{(\nu,E_{n+1})^{3}}\right|$
(4.39) $\geq\left(\frac{d}{\lambda_{n}\rho}-\left|\frac{\beta b_{n}}{(\nu,E_{n+1})}+\frac{2\alpha(e_{n},E_{n+1})}{(\nu,E_{n+1})^{3}}\right|\right)\frac{CdF^{2}\log h_{11}}{\lambda_{n}\lambda_{n-1}\rho}$
From the inequality (2.2) and conditions of case 3.2, we have
$\frac{1}{\lambda_{n}}\geq\frac{1}{C(n)F}\cdot\lambda_{n-1}\geq\frac{1}{C(n)F}\cdot\sqrt{\log{\lambda_{1}}}\quad;\quad\rho\leq\frac{1}{2}$
Then Consider the coefficient of inequality (4.39), we obtain
$\frac{d}{2\lambda_{n}\rho}-\left|\frac{\beta b_{n}}{(\nu,E_{n+1})}+\frac{2\alpha(e_{n},E_{n+1})}{(\nu,E_{n+1})^{3}}\right|\geq\frac{d\cdot\sqrt{\log{\lambda_{1}}}}{C(n)F}-\left|\frac{\beta b_{n}}{(\nu,E_{n+1})}+\frac{2\alpha(e_{n},E_{n+1})}{(\nu,E_{n+1})^{3}}\right|.$
This is clearly non-negative since we assume $\lambda_{1}$ is sufficiently large.
Hence by the discussion above, we obtain:
(4.40) $\frac{F^{nn}h_{11n}^{2}}{20h_{11}\log h_{11}}\geq\frac{d}{2\lambda_{n}\rho}\cdot\frac{CdF^{2}\log h_{11}}{\lambda_{n}\lambda_{n-1}\rho}\geq\frac{Cd^{2}F\log h_{11}}{\lambda_{n}\rho^{2}}.$
Substituting inequality (4.40) into (4.17), we derive:
$F^{ij}P_{ij}$ $\geq\frac{4\sum_{i}F^{ii}}{\rho}-\frac{16F^{ii}a_{i}^{2}}{\rho^{2}}-C\left(\frac{1}{\rho}+\beta+\alpha\right)+\frac{F^{ii}h_{11i}^{2}}{20h_{11}\log h_{11}}$
$\quad+\frac{Fh_{11}}{\log h_{11}}+\alpha\frac{F^{ii}h_{ii}^{2}}{(\nu,E_{n+1})^{2}}$
$\geq-\frac{C}{\rho\lambda_{n}}-\frac{C}{\rho^{2}\lambda_{n}}-C\left(\frac{1}{\rho}+\beta+\alpha\right)+\frac{Cd^{2}F\log h_{11}}{\lambda_{n}\rho^{2}}$
(4.41) $\quad+\frac{Fh_{11}}{\log h_{11}}+\alpha\frac{F^{ii}h_{ii}^{2}}{(\nu,E_{n+1})^{2}}.$
Since $h_{11}$ is sufficiently large, inequality (4) becomes
(4.42) $F^{ij}P_{ij}\geq-C\left(1+\beta+\alpha\right)+\frac{Fh_{11}}{\log h_{11}},$
which provides the estimate for this case. And this completes the proof.
Acknowledgements
This work was carried out while the author was a Master’s student at McMaster University, and constitutes the author’s Master’s thesis. The current version also corrects certain errors present in the original thesis manuscript. The author would like to thank Professor Siyuan Lu for helpful discussions and for bringing many related works to our attention. He is also deeply grateful to Bin Wang for numerous important discussions and for his careful suggestions regarding the correctness of this work. In particular, the author sincerely thanks Professor Guohuan Qiu for kindly reviewing the manuscript with great care.